OpenStudy (anonymous):

FOR A MEDAL Identify the slope and the Y intercept of y = -1/3x + 3 FOR A MEDAL

OpenStudy (anonymous):

y = mx + b y = -1/3x + 3 slope = m y - intercept = 3

OpenStudy (anonymous):

slope: \[-\frac{ 1 }{ 3 }\] y - intercept:

OpenStudy (anonymous):

y intercept: \[3\]

OpenStudy (anonymous):

thanks homie!

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

okaii, so since the slope-intercept form is y = mx + b, we can use this a=to graph the line. First we plot the y-intercept; which in this case is positive 3. Then we use the slope -1/3 and use rise over run. We use the rise over run from the y-intercept, count up or down for the rise, over right or left for the run, and put the next point. So, it would be up or down -1 on the y-axis and run 3 to the right or left. Then connect the two points and this is your line.